The daily milk production of a herd of cows is assumed to be Normally distributed with a mean of 39 liters, and standard deviation of 12.9 liters. A) On what proportion of days is daily production less than 12.5 liters? Answer= (Round your answer to 3 decimal places.) B) On what proportion of days is production more than 23.6 liters? Answer= (Round your answer to 3 decimal places.)
Answer #1

Let X denotes the daily milk production of a herd of cows.

X ~ Normal(39, 12.92)

a) The proportion of days is daily production less than 12.5 liters

=P(X<12.5) = P(\frac{X-39}{12.9}<\frac{12.5-39}{12.9})

= P(Z <\frac{12.5-39}{12.9})

= P(Z <-2.05426)

= \Phi(-2.05426)

= 0.0199751

b) The proportion of days is production more than 23.6 liters

=P(X>23.6) =1-P(X\leq 23.6) =1- P(\frac{X-39}{12.9}\leq \frac{23.6-39}{12.9})

=1- P(Z \leq \frac{23.6-39}{12.9})

=1- P(Z \leq -1.19380)

=1- \Phi(-1.19380)

=1-0.116278

=0.883722